Goods costing $18.16 at the beginning of the year 1957 cost $174.18 at the beginning of the year 2002. Find the average annual inflation rate over this time period. Express your answer the nearest thousandth of a percent.
Answer:_______ %

Answer by jorel555(1290)

2002-1957=45
17418/1816=9.5914
9.5914/45 years= annual inflation rate of 21.3142% ☺☺☺